OpenStudy (anonymous):

[4_06]Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

use the "point slope" formula \[y-y_1=m(x-x_1)\] replace \(x_1\) by \(-3\) and \(y_1\) by \(-1\) and \(m\) by \(4\)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ (&3\quad ,&-1) \end{array} \\\quad \\ slope = m= 4 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (jdoe0001):

hmmm \(\bf \begin{array}{lllll} &x_1&y_1\\ &(-3\quad ,&-1) \end{array} \\\quad \\ slope = m= 4 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

y = -2x - 7 y = -2x +5 y = 2x + 5 y = 2x - 7 here are my choices

OpenStudy (anonymous):

@satellite73 then what? i got y-1=4(x-3)

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(-3\quad ,&-1) \end{array} \\\quad \\ slope = m= 4 \\ \quad \\ y-y_1=m(x-x_1)\implies y-(-1)=4(x-(-3)\)

OpenStudy (jdoe0001):

then just solve for "y" to get the equation

OpenStudy (anonymous):

@jdoe0001 can you jts give the answer?

OpenStudy (jdoe0001):

makes us wonder why you weren't given the answer sheet along with the exercise do you know why?

OpenStudy (anonymous):

ok , i understand. but i donr get it. i tried solving for y and got something ridiculous

OpenStudy (jdoe0001):

well... tis ok... what did you get?

OpenStudy (anonymous):

y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

ohhhhhh , it is the first one?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmmm.... those look like choices, rather than your simplification

OpenStudy (anonymous):

i didnt have to simplfy , these are my choices?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(-3\quad ,&-1) \end{array} \\\quad \\ slope = m= 4 \\ \quad \\ y-y_1=m(x-x_1)\implies y-(\color{red}{-1})=4(x-(\color{red}{-3}))\)

OpenStudy (jdoe0001):

recall that - times - = +